Suppose a marketing company wants to determine the current proportion of customers who click on ads on their smartphones. It was estimated that the current proportion of customers who click on ads on their smartphones is 0.68. How many customers should the company survey in order to be 97% confident that the margin of error is 0.23 for the confidence interval of true proportion of customers who click on ads on their smartphones?
Answer #1

Solution :

Given that,

\hat p = 0.68

1 - \hat p = 0.32

margin of error = E = 0.23

Z\alpha/2 = 2.17

sample size = n = (Z\alpha / 2 / E)2 * \hat p * (1 - \hat p )

= (2.17 / 0.23)2 * 0.68 * 0.32

= 20

20 customers